# Training-Free Text-Guided Color Editing with Multi-Modal Diffusion Transformer

### v2

@@
Abstract.
Text-guided color editing in images and videos is a fundamental yet unsolved problem, requiring fine-grained manipulation of color attributes, including albedo, light source color, and ambient lighting, while preserving physical consistency in geometry, material properties, and light-matter interactions. Existing training-free methods offer broad applicability across editing tasks but struggle with precise color control and often introduce visual inconsistency in both edited and non-edited regions. In this work, we present ColorCtrl, a training-free color editing method that leverages the attention mechanisms of modern Multi-Modal Diffusion Transformers (MM-DiT). By disentangling structure and color through targeted manipulation of attention maps and value tokens, our method enables accurate and consistent color editing, along with word-level control of attribute intensity. Our method modifies only the intended regions specified by the prompt, leaving unrelated areas untouched. Extensive experiments on both SD3 and FLUX.1-dev demonstrate that ColorCtrl outperforms existing training-free approaches and achieves state-of-the-art performances in both edit quality and consistency. Furthermore, our method surpasses strong commercial models such as FLUX.1 Kontext Max and GPT-4o Image Generation in terms of consistency. When extended to video models like CogVideoX, our approach exhibits greater advantages, particularly in maintaining temporal coherence and editing stability. Finally, our method also generalizes to instruction-based editing diffusion models such as Step1X-Edit and FLUX.1 Kontext dev, further demonstrating its versatility. Here is the project website.
copyright: nonecopyright: none
Figure 1. Text-conditioned color editing. Our method, ColorCtrl with FLUX.1-dev, edits colors across multiple materials while preserving light-matter interactions. For example, in the fourth case, the ball’s color, its water reflection, specular highlights, and even small droplets on the glass have all been changed. It also enables fine-grained control over the intensity of specific descriptive terms.
1. Introduction
@@
Traditional image processing methods have been widely commercialized in professional software such as Photoshop, serving billions of users worldwide. However, the steep learning curve and significant manual effort required make it difficult to achieve widespread accessibility. Moreover, such tools are not well-suited for automated batch processing and present challenges when applied to video-related tasks. Recently, diffusion models have demonstrated remarkable capabilities in generating high-quality images that adhere to physical principles of color and illumination. This has spurred growing interest in leveraging their generative power to tackle the above challenges, with controllability emerging as a critical factor. Although many methods (Magar et al., 2025; Zhang et al., 2025) fine-tune diffusion models for controllable editing, they usually require large-scale datasets and complex training pipelines, and are often constrained to narrow domains or specific edit types. On the other hand, training-free methods (Jiao et al., 2025) have gained popularity for a wide range of image editing tasks due to their generality and ease of use. Despite their success in many scenarios, they still struggle with fine-grained color control and often introduce inconsistencies in edited and non-edited regions, making precise color editing an unresolved challenge.
The recent architectural shift in diffusion models from U-Net (Rombach et al., 2022) to the Multi-Modal Diffusion Transformer (MM-DiT) (Esser et al., 2024) opens up new opportunities for training-free color editing, for two main reasons: (1) the new architecture allows for scaling up both data and model capacity, enabling better adherence to physical priors; and (2) the improved fusion of text and vision modalities supports more flexible and precise attention-control editing strategies.
@@
To verify the effectiveness of our method, we conduct extensive experiments showing that ColorCtrl outperforms existing training-free approaches on both Stable Diffusion 3 Medium (a.k.a., SD3) (Esser et al., 2024) and FLUX.1-dev (Labs, 2024). To better understand the gap between open-source and commercial models, we further compare ColorCtrl with two strong commercial baselines: FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (a.k.a., GPT-4o) (OpenAI, 2025). ColorCtrl achieves more natural color editing and substantially better preservation of visual consistency, including background and structural fidelity. In addition, our method is model-agnostic and can be seamlessly extended to video models such as CogVideoX (Yang et al., 2024), where the performance gap becomes even more pronounced. Finally, ColorCtrl can be integrated into instruction-based editing models such as Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025), demonstrating strong compatibility.
@@
2. Related Work
Text-to-image and video generation. Diffusion models with a U-Net backbone (Ho et al., 2020; Rombach et al., 2022; Guo et al., 2024) have largely replaced early GAN systems (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) due to superior image fidelity. However, the U-Net design scales poorly, leading to the adoption of Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among these, MM-DiT (Esser et al., 2024) has emerged as a widely adopted backbone for recent state-of-the-art models (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025b; Black Forest Labs, 2025), such as SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, as well as CogVideoX (Yang et al., 2024) for video generation. In this work, we propose an attention-control scheme that plugs into any MM-DiT model.
Text-guided editing. Training-free text-guided editing methods based on pre-trained diffusion models offer flexibility and efficiency. Existing approaches fall into two groups: (i) sampling-based methods, which steer generation by injecting controlled noise or inversion (Jiao et al., 2025; Huberman-Spiegelglas et al., 2024; Xu et al., 2023; Kulikov et al., 2024; Yan et al., 2025); (ii) attention-based methods, which modify attention maps, starting with Prompt-to-Prompt (Hertz et al., 2023) and its image/video variants (Chen et al., 2024; Wang et al., 2025; Liu et al., 2024; Cao et al., 2023; Cai et al., 2025; Rout et al., 2025; Xu et al., 2025). Among these, DiTCtrl is the only method exploring attention control in MM-DiT, but ColorCtrl differs in key ways: DiTCtrl only applies mask extraction during long video generation, not color editing, while our improved method is more robust. Additionally, DiTCtrl’s attribute re-weighting disrupts attention consistency, causing inconsistent geometries, while our method avoids this. Furthermore, ColorCtrl works across all layers and timesteps, unlike DiTCtrl, which requires careful layer selection. Different from DiTCtrl, we operate in attention maps rather than key and value tokens. More recently, methods (Liu et al., 2025a; Brooks et al., 2023) like FLUX.1 Kontext (Black Forest Labs, 2025) and GPT-4o (OpenAI, 2025) have shown convenient creative workflows in training with synthetic instruction-response pairs for fine-tuning a diffusion model for image editing.
@@ 3. Method
3.1. Task Formulation: Text-guided Color Editing
Rendering process. To provide a clearer and more precise description of the conditions required for film industry-level color changing, we define the rendering process of a frame by
(1) $\textbf{I}\;=\;\mathcal{R}\bigl{(}G,\;L,\;A,\;S,\;C\bigr{)},$
@@ 3.1. Task Formulation: Text-guided Color Editing
Editing task. According to the predefined image rendering process in Eq. 1, we formulate the consistent editing task in this work as follows. Given a source image I and text prompt pairs $q$ before and after editing [^1] An example of $q$: “white fox.” $\to$ “orange fox.” (source and target prompt respectively). that specifies which objects or lights to recolor and the desired target color, we aim to learn the following editing function $f(\cdot)$
(2) $f:(\textbf{I},q)\,\longmapsto\,\hat{\textbf{I}}=\mathcal{R}\bigl{(}G,\hat{L},\hat{A},S,C\bigr{)},$
@@
Figure 2. Pipeline of ColorCtrl. (a) Visualizes the attention mechanism in MM-DiT blocks. (b) Enables effective color editing while maintaining structural consistency. (c) Preserves colors in non-editing regions. (d) Applies attribute re-weighting to specific word tokens. Symbols in the source branch have no superscript. Symbols with a superscript indicate the target branch, and hats (e.g., $\hat{V}$, $\hat{M}$) denote edited outputs.
3.2. Preliminaries: Attention in MM-DiT Blocks
@@ 3.3. Structure Preservation
Following (Hertz et al., 2023; Cao et al., 2023), we divide the editing process into a source branch and a target branch. The source branch follows the original generation process, producing a source image and storing intermediate attention outputs. The target branch reuses these stored variables to generate edited results. Starting from a fixed random seed, the model can generate the desired object without applying any editing method. However, the resulting image often diverges significantly from the source, making it unsuitable for meaningful editing tasks that require structural consistency. To address this, all subsequent modifications are applied exclusively on the target branch, aiming to improve consistency without disrupting the intended edit. The process of keeping $G$, $S$, $C$, light positions, and scalar intensities fixed (as in constraints (C1)-(C3)) is referred to as structure preservation. We observe that the vision-to-vision part of the attention map inherently encodes rich knowledge about the parts of the scene that must remain unchanged. Given a source attention map $\bm{M}$, its vision-to-vision part is transferred from $\bm{M}$ to the corresponding part in the target attention map $\bm{M^{*}}$, producing an edited attention map $\bm{\hat{M}}$ that fully respects the structure preservation constraints, as described in Fig. 2 (b).
3.4. Color Preservation
@@
Table 1. Quantitative image results compared with training-free methods on PIE-Bench. Results for FireFlow on SD3 are omitted due to their consistency being worse than that obtained using fixed seeds.
Figure 3. Top row: SD3 results; bottom row: FLUX.1-dev results. (a) The edit prompt is “white fox” $\to$ “orange fox”. Left to right: source image, our full method, without color preservation, with swapped text-to-text part in structure preservation, and with swapped $V^{\text{text}}$ in color preservation. (b) The generation prompt is “a white fox in a forest”, and the token for mask extraction is “fox”. From left to right: the mask extracted from vision-to-text parts, and from text-to-vision parts.
Despite enforcing structure preservation, we observe that undesired changes such as color shifts can still occur in regions unrelated to the edit, as shown in Fig. 3 (a). To further localize the edit and reduce inconsistencies, edits should be confined to the intended regions, while preserving all other areas. Motivated by the approach of Cai et al. (2025), we first extract a binary mask $\bm{m}$ from vision-to-text parts of attention maps with a threshold $\epsilon$, which indicates the target editing region. In contrast to Cai et al. (2025), which averages the vision-to-text and text-to-vision parts to obtain the mask, we use only the vision-to-text parts, as they provide superior spatial localization for the target object, unlike the text-to-vision parts (see Fig. 3 (b)). Based on $\bm{m}$, we copy the value tokens from the non-editing regions in the vision part of the source $\bm{V}$ into the corresponding regions of the target $\bm{V^{*}}$, yielding the final value tokens $\bm{\hat{V}}$, as shwon in Fig. 2 (c). We refer to this procedure as color preservation. As shown in Fig. 3 (a), including the text part of value tokens $\bm{V^{\text{text}}}$ during value exchange significantly weakens text guidance in FLUX.1-dev, and leads to severe artifacts in SD3.
3.5. Attribute Re-Weighting
So far, our method enables powerful color editing capabilities. However, the granularity of control via text prompts remains limited. For example, if the user specifies a color like “dark yellow”, there is no way to explicitly control the degree of darkness. Existing re-weighting techniques in U-Net-based models primarily follow two approaches: (1) Scaling the text embedding of specific tokens before the diffusion process (Perry, 2023). However, this method is designed for CLIP-based (Radford et al., 2021) text encoders and is incompatible with MM-DiT-based models, which typically use T5 (Raffel et al., 2020). (2) Scaling the attention scores of specific tokens in the cross-attention map (Hertz et al., 2023). This is also inapplicable to MM-DiT, which relies solely on self-attention. Moreover, both approaches fail to maintain structural consistency during scaling, limiting their utility in high-fidelity editing tasks.
4. Experiments
4.1. Setup
Baselines. We compare our method against several training-free approaches built upon MM-DiT, including FireFlow (Deng et al., 2025), RF-Solver (Wang et al., 2025), SDEdit (Meng et al., 2022), DiTCtrl (Cai et al., 2025), FlowEdit (Kulikov et al., 2024), and UniEdit-Flow (Jiao et al., 2025). We focus exclusively on MM-DiT-based baselines, as prior work (Jiao et al., 2025; Deng et al., 2025) has shown that U-Net-based methods perform significantly worse. Accordingly, we exclude methods that cannot be adapted to the MM-DiT architecture. For commercial baselines, we compare with FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025).
@@ 4.1. Setup
Implementation. We conduct experiments on SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, and on CogVideoX-2B (Yang et al., 2024) for video generation. For instruction-based image editing, we use Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025). For FLUX.1-dev, Step1X-Edit, and FLUX.1 Kontext dev, we apply attention control to the single-stream attention layers, following Deng et al. (2025). Unless otherwise noted, we use the Euler sampler and adopt UniEdit-Flow (Jiao et al., 2025) for image inversion. Maintaining a balance between fidelity to the original image and the strength of the applied edits is a well-known trade-off in generative editing. To ensure a fair comparison across methods, we carefully tune the hyperparameters for each baseline. Additional details are provided in Appendix A.
Figure 4. Qualitative image results compared with training-free methods and commercial models on PIE-Bench. The top three rows are generated using FLUX.1-dev, while the bottom two are generated using SD3. Best viewed with zoom-in.

### v3

@@
Abstract
Text-guided color editing in images and videos is a fundamental yet unsolved problem, requiring fine-grained manipulation of color attributes, including albedo, light source color, and ambient lighting, while preserving physical consistency in geometry, material properties, and light-matter interactions. Existing training-free approaches provide broad applicability across editing tasks but struggle with precise color control and often introduce visual inconsistency in both edited and non-edited regions. In this work, we present ColorCtrl, a training-free color editing method that leverages the attention mechanisms of modern Multi-Modal Diffusion Transformers (MM-DiT). By disentangling structure and color through targeted manipulation of attention maps and value tokens, our method enables accurate and consistent color editing, along with word-level control of attribute intensity. Our method modifies only the intended regions specified by the prompt, leaving unrelated areas untouched. Extensive experiments on both SD3 and FLUX.1-dev demonstrate that ColorCtrl outperforms existing training-free approaches and achieves state-of-the-art performances in both edit quality and consistency. Furthermore, our method surpasses strong commercial models such as FLUX.1 Kontext Max and GPT-4o Image Generation in terms of consistency. When extended to video models like CogVideoX, our approach exhibits greater advantages, particularly in maintaining temporal coherence and editing stability. Finally, our method generalizes to instruction-based editing diffusion models such as Step1X-Edit and FLUX.1 Kontext dev, further demonstrating its versatility. Here is the website.
1 Introduction
@@
Traditional image processing methods have been widely commercialized in professional software such as Photoshop, serving billions of users worldwide. However, the steep learning curve and significant manual effort required make it difficult to achieve widespread accessibility. Moreover, such tools are not well-suited for automated batch processing and cause problems when applied to video-related tasks. Recently, diffusion models have demonstrated remarkable capabilities in generating high-quality images that adhere to physical principles of color and illumination. This has spurred growing interest in leveraging their generative power to address the above challenges, with controllability emerging as a critical factor. Although many methods (Magar et al., 2025; Zhang et al., 2025) fine-tune diffusion models for controllable editing, they usually require large-scale datasets and complex training pipelines, and are often constrained to narrow domains or specific edit types. On the other hand, training-free methods (Jiao et al., 2025) have gained popularity for a wide range of image editing tasks due to their generality and ease of use. Despite their success in many scenarios, they still struggle with fine-grained color control and often introduce inconsistencies in edited and non-edited regions, making precise color editing an unresolved challenge.
The recent architectural shift in diffusion models from U-Net (Rombach et al., 2022) to the Multi-Modal Diffusion Transformer (MM-DiT) (Esser et al., 2024) brings new opportunities for training-free color editing, for two main reasons: (1) the new architecture allows for scaling up both data and model capacity, enabling better adherence to physical priors; and (2) the improved fusion of text and vision modalities supports more flexible and precise attention-control editing strategies.
Figure 1: Text-conditioned color editing. Our method, ColorCtrl with FLUX.1-dev, edits colors across multiple materials while preserving light-matter interactions. For example, in the fourth case, the ball’s color, its water reflection, specular highlights, and even small droplets on the glass have all been changed. It also enables fine-grained control over the intensity of specific descriptive terms.
@@
To verify the effectiveness of our method, we conduct extensive experiments showing that ColorCtrl outperforms existing training-free approaches on both Stable Diffusion 3 Medium (a.k.a., SD3) (Esser et al., 2024) and FLUX.1-dev (Labs, 2024). To better understand the gap between open-source and commercial models, we further compare ColorCtrl with two strong commercial baselines: FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (a.k.a., GPT-4o) (OpenAI, 2025). ColorCtrl achieves more natural color editing and substantially better preservation of visual consistency, including background and structural fidelity. In addition, our method is model-agnostic and can be seamlessly extended to video models such as CogVideoX (Yang et al., 2024), where the performance gap becomes even more pronounced. Finally, ColorCtrl can be integrated into instruction-based editing models such as Step1X-Edit (Liu et al., 2025b) and FLUX.1 Kontext dev (Black Forest Labs, 2025), demonstrating strong compatibility.
@@
2 Related Work
Text-to-image and video generation.
Diffusion models with a U-Net backbone (Ho et al., 2020; Rombach et al., 2022; Guo et al., 2024) have largely replaced early GAN systems (Reed et al., 2016; Yu et al., 2023; Wang et al., 2023) due to superior image fidelity. However, the U-Net design scales poorly, leading to the adoption of Diffusion Transformers (DiT) (Peebles and Xie, 2023). Among these, MM-DiT (Esser et al., 2024) has emerged as a widely adopted backbone for recent state-of-the-art models (Esser et al., 2024; AI, 2024; Labs, 2024; Yang et al., 2024; Kong et al., 2024; Liu et al., 2025a; Ma et al., 2025; Black Forest Labs, 2025), such as SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, as well as CogVideoX (Yang et al., 2024) for video generation. In this work, we propose an attention control that plugs into any MM-DiT model.
Text-guided editing.
Training-free text-guided editing methods based on pre-trained diffusion models offer flexibility and efficiency. Existing approaches fall into two groups: (i) sampling-based methods, which steer generation by injecting controlled noise or inversion (Jiao et al., 2025; Huberman-Spiegelglas et al., 2024; Xu et al., 2023; Kulikov et al., 2024; Yan et al., 2025); (ii) attention-based methods, which modify attention maps, starting with Prompt-to-Prompt (Hertz et al., 2023) and its image/video variants (Chen et al., 2024; Wang et al., 2025; Liu et al., 2024; Cao et al., 2023; Cai et al., 2025; Rout et al., 2025; Xu et al., 2025; Yin et al., 2025). Among these, DiTCtrl is the only method exploring attention control in MM-DiT, but ColorCtrl differs in key ways: DiTCtrl only applies mask extraction during long video generation, not editing, while our improved method is more robust. Additionally, DiTCtrl’s re-weighting disrupts attention consistency, causing inconsistent geometries, and TextCrafter (Du et al., 2025) exhibits a similar issue, while ours avoids this. Furthermore, ColorCtrl works across all layers and timesteps, unlike DiTCtrl, which requires careful layer selection. Different from DiTCtrl, we operate in attention maps rather than key and value tokens. More recently, methods (Liu et al., 2025b; Brooks et al., 2023) like FLUX.1 Kontext (Black Forest Labs, 2025) and GPT-4o (OpenAI, 2025) have shown convenient creative workflows in training with synthetic instruction-response pairs for fine-tuning a diffusion model for image editing.
@@ 3 Method
3.1 Task Formulation: Text-guided Color Editing
Rendering process.
To provide a clearer and more precise description of the conditions required for film industry-level color changing, we define the rendering process of a frame by
$\textbf{I}\;=\;\mathcal{R}\bigl(G,\;L,\;A,\;S,\;C\bigr),$ (1)
@@ 3.1 Task Formulation: Text-guided Color Editing
Editing task.
According to the predefined image rendering process in Eq. 1, we formulate the consistent editing task in this work as follows. Given a source image I and text prompt pairs $q$ before and after editing [^1] An example of $q$: “white fox.” $\to$ “orange fox.” (source and target prompt respectively). that specifies which objects or lights to recolor and the desired target color, we aim to learn the following editing function $f(\cdot)$
$f:(\textbf{I},q)\,\longmapsto\,\hat{\textbf{I}}=\mathcal{R}\bigl(G,\hat{L},\hat{A},S,C\bigr),$ (2)
@@
Figure 2: Pipeline of ColorCtrl. (a) Visualizes the attention mechanism in MM-DiT blocks. (b) Enables color editing while maintaining structural consistency. (c) Preserves colors in non-editing regions. (d) Applies attribute re-weighting to specific tokens. Symbols in the source branch have no superscript. Symbols with a superscript indicate the target, and hats (e.g., $\hat{V}$, $\hat{M}$) denote outputs.
3.2 Preliminaries: Attention in MM-DiT Blocks
@@ 3.3 Structure Preservation
Following Hertz et al. (2023); Cao et al. (2023), we divide the editing process into a source branch and a target branch. The source branch follows the original generation process, producing a source image and storing intermediate attention outputs. The target branch reuses these stored variables to generate edited results. Starting from a fixed random seed, the model can generate the desired object without applying any editing method. However, the resulting image often diverges significantly from the source, making it unsuitable for meaningful editing tasks that require structural consistency. To address this, all subsequent modifications are applied exclusively on the target branch, aiming to improve consistency without disrupting the intended edit. The process of keeping $G$, $S$, $C$, light positions, and scalar intensities fixed (as in constraints (C1)-(C3)) is referred to as structure preservation. We observe that the vision-to-vision part of the attention map inherently encodes rich knowledge about the parts of the scene that must remain unchanged. Given a source attention map $\bm{M}$, its vision-to-vision part is transferred from $\bm{M}$ to the corresponding part in the target attention map $\bm{M^{*}}$, producing an edited attention map $\bm{\hat{M}}$ that fully respects the structure preservation constraints, as described in Fig. 2 (b).
3.4 Color Preservation
Figure 3: Top row: SD3 results; bottom row: FLUX.1-dev results. (a) The edit prompt is “white fox” $\to$ “orange fox”. Left to right: source image, our full method, without color preservation, with swapped text-to-text part in structure preservation, and with swapped $V^{\text{text}}$ in color preservation. (b) The generation prompt is “a white fox in a forest”, and the token for mask extraction is “fox”. From left to right: the mask extracted from vision-to-text parts, and from text-to-vision parts.
Despite enforcing structure preservation, we observe that undesired changes such as color shifts can still occur in regions unrelated to the edit, as shown in Fig. 3 (a). To further localize the edit and reduce inconsistencies, edits should be confined to the intended regions, while preserving all other areas. Motivated by the approach of Cai et al. (2025), we first extract a binary mask $\bm{m}$ from vision-to-text parts of attention maps with a threshold $\epsilon$, which indicates the target editing region. In contrast to Cai et al. (2025), which averages the vision-to-text and text-to-vision parts to obtain the mask, we use only the vision-to-text parts, as they provide superior spatial localization for the target, unlike the text-to-vision parts (see Fig. 3 (b)). Based on $\bm{m}$, we copy the value tokens from the non-editing regions in the vision part of the source $\bm{V}$ into the corresponding regions of the target $\bm{V^{*}}$, yielding the final value tokens $\bm{\hat{V}}$, as shwon in Fig. 2 (c). We refer to this procedure as color preservation. As shown in Fig. 3 (a), including the text part of value tokens $\bm{V^{\text{text}}}$ during value exchange significantly weakens text guidance in FLUX.1-dev, and leads to severe artifacts in SD3.
3.5 Attribute Re-Weighting
So far, our method enables powerful color editing capabilities. However, the granularity of control via text prompts remains limited. For example, if the user specifies a color like “dark yellow”, there is no way to explicitly control the degree of darkness. Existing re-weighting techniques in U-Net-based models primarily follow two approaches: (1) Scaling the text embedding of specific tokens before the diffusion process (Perry, 2023). However, this method is designed for CLIP-based (Radford et al., 2021) text encoders and is incompatible with MM-DiT-based models, which typically use T5 (Raffel et al., 2020). (2) Scaling the attention scores of specific tokens in the cross-attention map (Hertz et al., 2023). This is also inapplicable to MM-DiT, which relies solely on self-attention. Moreover, both existing approaches fail to maintain structural consistency during scaling, limiting their utility in high-fidelity editing tasks.
4 Experiments
4.1 Setup
Baselines.
We compare our method against several training-free approaches built upon MM-DiT, including FireFlow (Deng et al., 2025b), RF-Solver (Wang et al., 2025), SDEdit (Meng et al., 2022), DiTCtrl (Cai et al., 2025), FlowEdit (Kulikov et al., 2024), and UniEdit-Flow (Jiao et al., 2025). We focus exclusively on MM-DiT-based baselines, as prior work (Jiao et al., 2025; Deng et al., 2025b) has shown that U-Net-based methods perform significantly worse. Accordingly, we exclude methods that cannot be adapted to the MM-DiT architecture. For commercial baselines, we compare with FLUX.1 Kontext Max (Black Forest Labs, 2025) and GPT-4o Image Generation (OpenAI, 2025).
Table 1: Quantitative image results compared with training-free methods on PIE-Bench. Results for FireFlow on SD3 are omitted due to their consistency being worse than using fixed seeds.
@@ 4.1 Setup
Implementation.
@@
We conduct experiments on SD3 (Esser et al., 2024) and FLUX.1-dev (Labs, 2024) for image generation, and on CogVideoX-2B (Yang et al., 2024) for video generation. For instruction-based image editing, we use Step1X-Edit (Liu et al., 2025b) and FLUX.1 Kontext dev (Black Forest Labs, 2025). For FLUX.1-dev, Step1X-Edit, and FLUX.1 Kontext dev, we apply attention control to the single-stream attention layers, following Deng et al. (2025b). Unless otherwise noted, we use the Euler sampler and adopt UniEdit-Flow (Jiao et al., 2025) for image inversion. Maintaining a balance between fidelity to the original image and the strength of the applied edits is a well-known trade-off in generative editing. To ensure a fair comparison across methods, we carefully tune the hyperparameters for each baseline. Additional details are provided in Appendix A.
Benchmark.
While prior editing methods (Hertz et al., 2023; Cao et al., 2023; Cai et al., 2025) typically lack standardized benchmark evaluation, we adopt prompts from the Change Color task in PIE-Bench (Ju et al., 2024), which includes 40 editing pairs, to better showcase the capabilities of our method. Although our approach is fully compatible with inversion methods, we adopt a noise-to-image setting during benchmark evaluations to better isolate and evaluate editing performance, removing the influence of reconstruction and inversion. This also allows us to reuse the same set of benchmark prompts for video diffusion models, enabling a consistent and fair comparison across both image and video domains. To further scale up the evaluation, we introduce a new benchmark called ColorCtrl-Bench, consisting of 300 prompt pairs in the same style as PIE-Bench (see Sec. A.4 for details). For all baselines, we adopt a fixed sampler and identical random seeds to ensure that source images are consistent, enabling reliable comparison across methods.
Evaluation protocol.
Unlike the original PIE-Bench, which evaluates structural similarity using structural distance (Tumanyan et al., 2022), we adopt the Structural Similarity Index (SSIM) (Wang et al., 2004) computed on Canny edge maps (Canny, 1986), following the approach of Zhao et al. (2023), for more accurate assessment. To evaluate the preservation of non-edited regions (a.k.a., BG Preservation), we compute PSNR and SSIM exclusively on those regions, which are annotated using Grounded SAM 2 (Ren et al., 2024) with dilation. Semantic alignment of the edits is assessed using CLIP similarity (Radford et al., 2021), applied to both the entire image and the edited regions.
4.2 Comparison with Training-Free Methods (Images)
Table 2: Quantitative image results compared with commercial models on PIE-Bench.